Let fn be the nth Fibonacci number. Prove that, for n > 0 [Hint: use strong induction]:

fn = 1/√5 [((1+√5)/2)n - ((1-√5)/2)n ]


1
Expert's answer
2021-10-26T17:58:35-0400

for n=1:

"F_1=1"


let for n=k:

"F_k=\frac{(\frac{1+\sqrt5}{2})^k-(\frac{1-\sqrt5}{2})^k}{\sqrt5}"


then for n=k+1:

"F_{k+1}=F_k+F_{k-1}=\frac{(\frac{1+\sqrt5}{2})^{k}-(\frac{1-\sqrt5}{2})^{k}}{\sqrt5}+\frac{(\frac{1+\sqrt5}{2})^{k-1}-(\frac{1-\sqrt5}{2})^{k-1}}{\sqrt5}="


"=\frac{1}{\sqrt 5}((\frac{1+\sqrt 5}{2})^k(1+\frac{2}{1+\sqrt 5})-(\frac{1-\sqrt 5}{2})^k(1+\frac{2}{1-\sqrt 5}))"


"(1+\frac{2}{1+\sqrt 5})(\frac{1-\sqrt 5}{1-\sqrt 5})=\frac{1+\sqrt 5}{2}"


"(1+\frac{2}{1-\sqrt 5})(\frac{1+\sqrt 5}{1+\sqrt 5})=\frac{1-\sqrt 5}{2}"


So:


"F_{k+1}=\frac{1}{\sqrt 5}((\frac{1+\sqrt 5}{2})^k(\frac{1+\sqrt 5}{2})-(\frac{1-\sqrt 5}{2})^k(\frac{1-\sqrt 5}{2}))=\frac{(\frac{1+\sqrt5}{2})^{k+1}-(\frac{1-\sqrt5}{2})^{k+1}}{\sqrt5}"